# Bounds on detection of Bell correlations with entangled ultra-cold atoms in optical lattices under occupation defects

### v1

@@ Abstract
Bell non-locality stems from quantum correlations effectively identified using inequalities. Spin chains, simulated with ultra-cold atoms in optical lattices, Rydberg atoms in tweezer arrays, trapped ions, or molecules, allow single-spin control and measurement. Therefore, they are suitable for studying fundamental aspects of these correlations and non-locality. Occupation defects, such as vacancies or multiple atoms occupying a single site due to imperfect system preparation, limit the detection of Bell correlations. We investigate their impact using a simplified toy model parameterized by the probability of a site being singly occupied. We derive the corresponding Bell inequality and identify the smallest probability that establishes a lower bound for detecting Bell correlations. We relate the bound to two physical parameters leading to defects in occupations: non-zero temperature and filling factor, focusing on entangled ultra-cold atoms in optical lattices. Finally, we numerically validate the predictions of the toy model by full many-body simulations.
@@ I Introduction
Quantum mechanics introduced groundbreaking concepts of non-local correlations and entanglement that challenged well-established principles of classical physics, including realism, causality and locality Einstein et al. (1935). A classical picture can be recovered at the price of introducing local hidden variables, and the corresponding (LV) theories are shown to satisfy Bell inequalities Bell (1964). Quantum correlations that violate Bell inequalities are inconsistent with LV theory and are, therefore, referred to as non-local Brunner et al. (2014a).
@@
The most robust method for entanglement certification is provided by violating Bell’s inequalities, as it avoids assumptions about the physical nature and degrees of freedom to be measured or the calibration of measurements Bowles et al. (2018). This is known as the device-independent scenario, which is a powerful resource in many quantum information tasks, such as self-testing Mayers and Yao (2004), randomness amplification and expansion Colbeck and Renner (2012); Brandão et al. (2016), quantum key distribution Acín et al. (2007); Roy-Deloison et al. (2024); Nadlinger et al. (2022), and quantum sensing and metrology Fröwis et al. (2019); Niezgoda and Chwedeńczuk (2021). Recent experimental demonstrations have employed various platforms, including entangled photons Aspect et al. (1982); Weihs et al. (1998); Giustina et al. (2015), spins in nitrogen-vacancy centers Hensen et al. (2015), superconducting circuits Storz et al. (2023), pairs of Josephson phase qubits Ansmann et al. (2009), and neutral and ultra-cold atoms Rosenfeld et al. (2017); Schmied et al. (2016). Many studies have concentrated on few-particle scenarios; however, correlations also naturally emerge in quantum many-body systems Tura et al. (2014).
@@
Atom assembly in optical lattices and optical tweezer arrays Yamamoto et al. (2016); Schine et al. (2022); Madjarov et al. (2019), as well as trapped ions and molecules Micheli et al. (2006); Wang et al. (2022), offer excellent approaches for studying many-body entanglement, with capabilities for single-atom preparation, control, and detection. In the above-mentioned platforms, entanglement can be generated dynamically from the initial coherent state well approximated by the one-axis twisting (OAT) model Kitagawa and Ueda (1993) where spin-squeezed and GHZ states are produced. Given the control and local measurements possible in such architectures and their scalability, the systems offer the powerful setup to generate and study fundamental aspects of Bell correlations and non-locality, enabling the exploration of quantum information concepts. Imperfections, however, e.g. in system preparation, may introduce occupation defects, limiting Bell inequalities’ application and entanglement certification.
@@
In this paper, we study the role of occupation defects in detecting Bell correlations focusing on $N$ two-level bosonic atoms in a spin-entangled state distributed among $M$ sites of an optical lattice, such that $N\leq M$. In the Bell scenario, we consider the measurement of two local collective spin observables at each lattice site where 00, $1$ or $2$ atoms are present. For each measurement result there are three possible outcomes: ±$\pm 1/2$ when the site is singly occupied, and 00 whenever a defect such as an empty or doubly occupied site is present, as illustrated in Fig. 1. We introduce a simplified toy model where the atoms’ internal and external degrees of freedom are decoupled. The dynamics of the internal spin degrees of freedom, including possible entanglement among the spins, is assumed to be given by the OAT model with $N$ atoms. The external degrees of freedom are parameterized by the probability $p$ of having a single occupied site allowing vacant, double, etc. occupancy sites when $p<1$. We derive the corresponding $p$-parameterized Bell inequalities based on $M$- and two-sites Bell correlations Płodzień et al. (2022); Müller-Rigat et al. (2021). We analytically estimate the lowest (critical) value of the probability $p$ for violation of Bell inequality. We obtain $p_{c}\approx 1/\sqrt{2}$ for the $M$-site Bell correlations. For the two-site Bell correlations, we obtain $p_{c}=4/5$ and $p_{c}=\sqrt{3}/2$ when $N=pM<M$ and $N=M$, respectively. These results are general and can be relevant for any platform where individual addressing of spins is possible  Richerme et al. (2014); Kajtoch et al. (2018a); Płodzień et al. (2020a); He et al. (2019); Comparin et al. (2022a, b); Mamaev et al. (2023); Pelayo et al. (2023), for example as in recent experiments using an array of trapped ions Franke et al. (2023) and Rydberg atoms Bornet et al. (2023); Eckner et al. (2023) demonstrating generation of entanglement in terms of scalable spin-squeezing with tens of spins.
@@
Figure 1: (a) A chain of $M$ spins is created by two-level ultra-cold atoms loaded into an optical lattice. Bell scenario considered: on each of $M$ subsystems (sites) two different measurement are performed $\hat{s}^{(j)}_{\alpha}$, with $\alpha=0,1$, locally leading to three measurement outcomes $r^{(j)}_{1},\,r^{(j)}_{2},r^{(j)}_{3}\in\{-1/2,+1/2,0\}$. The M-site Bell correlator, is tailored to detect GHZ states, as the one shown in (b). The two-site Bell correlator detects spin-squeezed states, as illustrated in (c).
@@
We explore the predictions of the toy model by performing full many-body numerical simulations of ultra-cold atoms in an optical lattice. We consider two distinct protocols for Bell correlations generation, which differ by the source of the occupation defects.
@@
In the first one, Bell correlations between individual atoms are generated in the Mott insulating regime via a weak inhomogeneous magnetic field or interaction anisotropy Hernández Yanes et al. (2024, 2022), and the considered occupation defects are vacant sites (holes). The probability $p$ of single occupancy sites is the filling factor $f=N/M$. Our numerical simulations confirm predictions of the toy model of the lower bound on the filling factor concerning the detection of the $M$-sites Bell correlations. However, the detection of the two-site Bell correlations in the lattice system is hindered whenever $f<1$ for fixed positions of holes. The movement of holes turned out to be crucial to lower the numerical bound on the filling factor, allowing for the detection of two-site Bell correlations up to $f\approx p_{c}$, as predicted by the toy model.
@@
In the second protocol, entanglement is generated in the superfluid regime using atom-atom interactions Kajtoch et al. (2018b); Płodzień et al. (2020a, 2022) and next is transferred to the Mott phase by increasing the lattice height. The main source of defects is imposed by non-zero temperature $T$ of an initial sample when $N=M$. We relate analytically the probability $p$ with the initial temperature $T$ of the system in the limit of an isoentropic transformation that leaves the system in thermal equilibrium at each moment. In this case, we determine analytically the upper bound for the initial temperature corresponding to $p_{c}$ below which $M$- and two-site Bell correlations can be detected.
The paper is organized as follows: In Secs. II and III, we provide the theoretical framework for the Bell scenario and toy model. This is applied to the system composed of ultra-cold atoms in an optical lattice presented in Sec. IV via analytical and numerical approaches. In Secs. V and VI we present the results for the first and second protocols, respectively, discussing the role of defects introduced by holes and non-zero initial temperature of the system. We conclude with further discussion and outlook in Sec. VII.
@@ II Bell Scenario
We consider $M$ spatially separated subsystems labelled $i=1,...,M$, sharing a quantum $M$-partite state described by the density operator $\hat{\rho}$. In each subsystem, $k$ observables labeled $\alpha=0,...,k-1$, can be measured, each having $d$ possible outcomes. A Bell experiment, as shown in Fig. 1-(a), involves choosing an arbitrary observable for each of the $M$ subsystems $\alpha=\{\alpha_{i}\}_{i=1}^{M}$, and recording the measurement results $r=\{r^{(i)}\}_{i=1}^{M}$. The goal is to determine the $M$-partite probability distribution $P_{M}(r|\alpha)$ for the outcomes $r$ given the settings $\alpha$ for all possible choices of settings $\alpha$. In the LV theory Brunner et al. (2014b), any probability distribution $P_{M}(r|\alpha)$ can be written as
@@
In the Bell experiment described above, the possibility of decomposing the measured probability distribution $P_{M}(r|\alpha)$ into the form (1) constitutes the locality condition for correlations between the $M$ subsystems under study. Conversely, if the measured probability distribution $P_{M}(r|\alpha)$ cannot be written in the form (1), the correlations present in the system are non-local. This decomposition corresponds to the most general locality condition, as defined by Bell. In practice, as we will see later, this condition is often formulated as a Bell inequality.
In this paper, we consider an optical lattice with $M$ sites, each containing zero, one or two two-level atoms. Each atom with two internal states $a$ and $b$ is an effective spin $1/2$, and we can express local collective spins in the quantized form as $\hat{s}_{x}^{(j)}=(\hat{s}_{+}^{(j)}+\hat{s}_{-}^{(j)})/2$, $\hat{s}_{y}^{(j)}=(\hat{s}_{+}^{(j)}-\hat{s}_{-}^{(j)})/(2i)$, $\hat{s}^{(j)}_{z}=(\hat{a}_{j}^{\dagger}\hat{a}_{j}-\hat{b}_{j}^{\dagger}\hat{b}_{j})/2$ with $\hat{s}_{+}^{(j)}=\hat{a}_{j}^{\dagger}\hat{b}_{j}$, $\hat{s}_{-}^{(j)}=(\hat{s}_{+}^{(j)})^{\dagger}$.
@@
as illustrated in Fig.1-(a), where the result $r=0$ is assigned to any measurement result different from ±$\pm 1/2$. The probability distribution $P_{M}(r|\alpha)$ to obtain the results $r$ given the settings $\alpha$ can be theoretically calculated, in terms of the density operator $\hat{\rho}$ describing the system’s state, as
@@
In the next Section, we derive two Bell inequalities using this framework. The first one, relying on $M$-site correlations, is mainly useful for highly entangled states with a small number of lattice sites (see, e.g., Fig. 1-(b)). The second one, relying on two-site correlations, is more applicable to spin-squeezed states with a large number of lattice sites (see, e.g., Fig. 1-(c)). For both cases, we introduce a simplified toy model accounting for occupation defects that result in measurements outcomes $r^{(j)}\neq\pm 1/2$.
@@ III Toy model
In the lattice system considered one starts from a product state $|x\rangle^{\otimes N}$ where each atom is in a coherent superposition of two internal states, i.e. coherent spin state (CSS) along the $x$ direction. The entanglement between spins is dynamically obtained using many-body interactions by different protocols (see e.g., Sec. V and Sec. VI) that can be effectively described, in some limit, by the OAT model
@@
where $\chi$ quantifies the strength of interactions and $\hat{S}_{\sigma}$ represents the collective spin operator of $N$ two-level atoms with $\sigma=x,y,z$. In (5) we have $\sigma=z$. The OAT model generates spin-squeezed states as well as non-Gaussian entangled states, including the GHZ state Ferrini et al. (2008).
In order to study the role of occupation defects in the detection of Bell correlations, we introduce an approximate toy model where the internal and external degrees of freedom of the atoms are decoupled
@@
$\hat{\rho}_{\mathrm{extern}}=\bigotimes_{j=1}^{M}\hat{\rho}_{j}\quad\textrm{with}\quad\hat{\rho}_{j}=p|1\rangle_{j}{}_{j}\langle 1|+(1-p)\hat{\rho}^{\perp}_{j},$ (8)
@@ III.1 $M$-site Bell correlations
and $\mathcal{E}_{M}$ is the $M$-site Bell correlator. We choose, in each site $j$, the settings $\alpha=0$ and $\alpha=1$ corresponding to $\hat{s}_{y}^{(j)}$ and $\hat{s}_{z}^{(j)}$ respectively Płodzień et al. (2022), to form
@@
When considering the toy model (6), the only nonzero contribution to ${\cal E}_{M}$ (15) comes from the sites occupied by a single spin. We have
@@
where $p^{M}$ represents the probability that the $M$ sites are occupied with a single atom. In the non-Gaussian regime of the OAT dynamics (7), a macroscopic superposition of coherent states is created at some particular instants $\chi\tau=\pi/q$ labelled by an even integer $q=2,4,6,…,M$, with $q=2$ for the GHZ state Płodzień et al. (2022); Ferrini et al. (2008). The $M$-site Bell correlator corresponding to these non-Gaussian spin states, for $p=1$, reads $\mathcal{E}_{M}^{(p=1)}\eqsim 1/q^{2}$, see Eq.(17) in Płodzień et al. (2022). By replacing in (16), one can derive a critical value of $p$ below which the $M$-site Bell correlations present in the states generated at $\chi\tau=\pi/q$ cannot be detected

### v2

@@ Abstract
Bell non-locality stems from quantum correlations effectively identified using inequalities. Spin chains, simulated with ultra-cold atoms in optical lattices, Rydberg atoms in tweezer arrays, trapped ions, or molecules, allow single-spin control and measurement. Therefore, they are suitable for studying fundamental aspects of these correlations and non-locality. Occupation defects, such as vacancies or multiple atoms occupying a single site due to imperfect system preparation, limit the detection of Bell correlations. We study their effects with the help of a simplified toy model parameterised by the probability $p$ of having a single occupation for a given site. Within this model, and for entangled systems obtained by one-axis twisting evolution from an initial factorised state, we derive two Bell inequalities, one based on many-site correlations and the other on two-site correlations, and identify the smallest probability $p$ that allows the Bell inequalities violation to be detected. We then consider two physical realizations using entangled ultra-cold atoms in optical lattices where the parameter $p$ is related to a non-unitary filling factor and non-zero temperature. We test the predictions of the toy model against exact numerical results.
@@ I Introduction
Quantum mechanics introduced groundbreaking concepts of non-local correlations and entanglement that challenged well-established principles of classical physics, including realism, causality and locality [1]. A classical picture can be recovered at the price of introducing local hidden variables, and the corresponding local hidden variable (LV) theories are shown to satisfy Bell inequalities [2]. Quantum correlations that violate Bell inequalities are inconsistent with LV theory and are, therefore, referred to as non-local [3].
@@
The most robust method for entanglement certification is provided by violating Bell’s inequalities, as it is independent of assumptions about the physical nature and degrees of freedom to be measured or the calibration of measurements [4]. This is known as the device-independent scenario, which is a powerful resource in many quantum information tasks, such as self-testing [5], randomness amplification and expansion [6, 7], quantum key distribution [8, 9, 10], and quantum sensing and metrology [11, 12]. Recent experimental studies of Bell correlations have employed various platforms, including entangled photons [13, 14, 15], spins in nitrogen-vacancy centers [16], superconducting circuits [17], pairs of Josephson phase qubits [18], and ultra-cold neutral atoms [19, 20]. Many studies have concentrated on few-particle scenarios; however, non-local correlations also naturally emerge in quantum many-body systems [21].
@@
Ensembles of atoms in optical lattices and optical tweezer arrays [22, 23, 24], as well as trapped ions and molecules [25, 26], are ideal platforms for studying many-body entanglement, with capabilities for single-atom preparation, control, and detection. Entanglement can be generated in these platforms by various methods, effectively implementing a coherent evolution that is well approximated by the one-axis twisting (OAT) model [27] where spin-squeezed and GHZ states are produced. Imperfections, however, e.g. in system preparation, may introduce occupation defects, limiting the possibility of detecting Bell’s inequalities violation and applying them for entanglement certification.
In this paper, we study the role of occupation defects in detecting Bell correlations focusing on $N$ two-level bosonic atoms in a spin-entangled state distributed among $M$ sites of an optical lattice, with $N\leq M$.
In Section II we present the Bell scenario in which we consider the measurement of two local collective spin observables at each lattice site where 00, $1$ or $2$ atoms are present. For each measurement result, there are three possible outcomes: ±$\pm 1/2$ when the site is singly occupied, and 00 whenever a defect such as an empty or doubly occupied site is present, as illustrated in Fig. 1.
@@
In Section III we introduce a simplified toy model where the atoms’ internal and external degrees of freedom are decoupled. The dynamics of the internal spin degrees of freedom, including possible entanglement among the spins, is assumed to be given by the OAT model with $N$ atoms. The state of the external degrees of freedom is parameterized by the probability $p$ of having a single occupied site allowing vacant, double, etc. occupancy sites when $p<1$. We derive the corresponding $p$-parameterized Bell inequalities based on $M$- and two-sites Bell correlations [28, 29]. We analytically estimate the lowest (critical) value of the probability $p$ for violation of Bell inequality. We obtain $p_{c}=2^{1/M}/\sqrt{2}$ for the $M$-site Bell correlations measured for a GHZ state. For the two-site Bell correlations measured for a spin-squeezed state, we obtain $p_{c}=4/5$ and $p_{c}=\sqrt{3}/2$ when $N=pM<M$ and $N=M$, respectively. These results are general and may be relevant for any platform where individual addressing of spins is possible  [30, 31, 32, 33, 34, 35, 36, 37], for example as in recent experiments using an array of trapped ions [38] and Rydberg atoms [39, 40] demonstrating generation of entanglement in terms of scalable spin-squeezing with tens of spins.
@@
Figure 1: A chain of $M$ spins is created by two-level ultra-cold atoms loaded into an optical lattice. Bell scenario considered: on each of $M$ subsystems (sites) two different measurement are performed $\hat{s}^{(j)}_{\alpha}$, with $\alpha=0,1$, locally leading to three measurement outcomes $r^{(j)}_{1},\,r^{(j)}_{2},r^{(j)}_{3}\in\{-1/2,+1/2,0\}$. The outcome “0” is associated to empty or double occupied sites. The M-site Bell correlator is tailored for GHZ states, while the two-site Bell correlator is tailored for spin-squeezed states.
@@
We explore the prediction of the toy model by performing full many-body simulations of ultra-cold atoms in optical lattices described by the Bose-Hubbard defined in Section IV. We consider two distinct protocols for Bell correlations generation, which differ by the source of the occupation defects.
@@
In the first protocol, discussed in Section V, Bell correlations between individual atoms are generated in the Mott insulating regime via a weak inhomogeneous magnetic field or interaction anisotropy [41, 42], and the considered occupation defects are vacant sites (holes). The probability $p$ of single occupancy of a given site is here equal to the filling factor $f=N/M$. Our numerical simulations confirm predictions of the toy model of the lower bound on the filling factor concerning the detection of the $M$-sites Bell correlations. Concerning the two-site Bell correlations, we find that their detection is hindered whenever $f<1$ for fixed positions of holes. On the contrary, in the case of weak interaction anisotropy, the movement of holes lowers the minimal filling factor allowing for the detection of two-site Bell correlations down to $f\approx p_{c}$, as predicted by the toy model.
@@
In the second protocol, investigated in Section VI, entanglement is generated in the superfluid regime using atom-atom interactions [43, 32, 28] and transferred to the Mott phase by increasing the lattice height. The main source of defects when $N=M$ is imposed by non-zero temperature $T$ of an initial sample. We relate analytically the probability $p$ with the initial temperature $T$ of the system in the limit of an isoentropic transformation that leaves the system in thermal equilibrium at each moment. In this case, we determine analytically the upper bound for the initial temperature corresponding to $p_{c}$ below which $M$- and two-site Bell correlations can be detected.
Our results show that for large systems the lower bounds on the probability of defects are weakly dependent on the number of particles demonstrating the possibility of witnessing Bell correlations in lattice systems within the range of typical experimental parameters, as concluded with further discussion in Sec. VII.
@@ II Bell Scenario
We consider $M$ spatially separated subsystems labelled $i=1,...,M$, sharing a quantum $M$-partite state described by the density operator $\hat{\rho}$. In each subsystem, $k$ observables labeled $\alpha=0,...,k-1$, can be measured, each having $d$ possible outcomes. A Bell experiment, as shown in Fig. 1, involves choosing an arbitrary observable for each of the $M$ subsystems $\alpha=\{\alpha_{i}\}_{i=1}^{M}$, and recording the measurement results $r=\{r^{(i)}\}_{i=1}^{M}$. The goal is to determine the $M$-partite probability distribution $P_{M}(r|\alpha)$ for the outcomes $r$ given the settings $\alpha$ for all possible choices of settings $\alpha$. In the LV theory [44], any probability distribution $P_{M}(r|\alpha)$ can be written as
@@
In the Bell experiment described above, the possibility of decomposing the measured probability distribution $P_{M}(r|\alpha)$ into the form (1) constitutes the locality condition for correlations between the $M$ subsystems under study. Conversely, if the measured probability distribution $P_{M}(r|\alpha)$ cannot be written in the form (1), the correlations present in the system are non-local. In practice, as we will see later, this condition is often formulated as a Bell inequality.
In this paper, we consider an optical lattice with $M$ sites, each containing zero, one or two two-level atoms. Each atom with two internal states $a$ and $b$ is an effective spin $1/2$, and we can express local collective spins in the second quantized form as $\hat{s}^{(j)}_{z}=(\hat{a}_{j}^{\dagger}\hat{a}_{j}-\hat{b}_{j}^{\dagger}\hat{b}_{j})/2$ and $\hat{s}_{x}^{(j)}=(\hat{s}_{+}^{(j)}+\hat{s}_{-}^{(j)})/2$, $\hat{s}_{y}^{(j)}=(\hat{s}_{+}^{(j)}-\hat{s}_{-}^{(j)})/(2i)$, with $\hat{s}_{+}^{(j)}=\hat{a}_{j}^{\dagger}\hat{b}_{j}$, $\hat{s}_{-}^{(j)}=(\hat{s}_{+}^{(j)})^{\dagger}$.
@@
as illustrated in Fig.1, where the result $r=0$ is assigned to any measurement result different from ±$\pm 1/2$. The probability distribution $P_{M}(r|\alpha)$ to obtain the results $r$ given the settings $\alpha$ can be theoretically calculated, in terms of the density operator $\hat{\rho}$ describing the system’s state, as
@@
In the next Section, we derive two Bell inequalities using this framework. The first one, relying on $M$-site correlations, is mainly useful for highly entangled states with a small number of lattice sites (see, e.g., Fig. 1). The second one, relying on two-site correlations, is more suitable to spin-squeezed states with a large number of lattice sites (see, e.g., Fig. 1). For both cases, we introduce a simplified toy model accounting for occupation defects that result in measurement outcomes $r^{(j)}=0$.
@@ III Toy model
In the lattice system considered one starts from a product state $|x\rangle^{\otimes N}$ where each atom is in a coherent superposition of two internal states, i.e. a coherent spin state (CSS) along the $x$ direction. The entanglement between spins is dynamically obtained using many-body interactions by different protocols (see Sec. V and Sec. VI) that can be effectively described, in some limit, by unitary evolution with the OAT Hamiltonian
@@
where $\chi$ quantifies the strength of interactions and $\hat{S}_{\sigma}$ represents the collective spin operator of $N$ two-level atoms with $\sigma=x,y,z$. The OAT model generates spin-squeezed states as well as non-Gaussian entangled states, including the GHZ state [45, 46].
To study the role of occupation defects in the detection of Bell correlations, we introduce an approximate toy model where the internal and external degrees of freedom of the atoms are decoupled
@@
$\hat{\rho}_{\mathrm{ext}}=\bigotimes_{j=1}^{M}\hat{\rho}_{j}\quad\textrm{with}\quad\hat{\rho}_{j}=p|1\rangle_{j}{}_{j}\langle 1|+(1-p)\hat{\rho}^{\perp}_{j},$ (8)
@@ III.1 $M$-site Bell correlations
and $\mathcal{E}_{M}$ is the $M$-site Bell correlator. We choose, in each site $j$, the settings $\alpha=0$ and $\alpha=1$ corresponding to $\hat{s}_{y}^{(j)}$ and $\hat{s}_{z}^{(j)}$ respectively [28], to form
@@
When considering the toy model (6), the only nonzero contribution to ${\cal E}_{M}$ (15) comes from the sites occupied by a single atom. We have
@@
where $p^{M}$ represents the probability that the $M$ sites are occupied with a single atom. In the non-Gaussian regime of the OAT dynamics (7), a macroscopic superposition of coherent states is created at some particular instants $\chi\tau=\pi/q$ labelled by an even integer $q=2,4,6,…,M$, with $q=2$ for the GHZ state [28, 46]. The $M$-site Bell correlator corresponding to these non-Gaussian spin states, for $p=1$, reads $\mathcal{E}_{M}^{(p=1)}\eqsim 1/q^{2}$, see Eq.(17) in [28]. By replacing in (16), one can derive a critical value of $p$ below which the $M$-site Bell correlations present in the states generated at $\chi\tau=\pi/q$ cannot be detected